OpenStudy (anonymous):

find the value of x so that 6/24=-2/x

OpenStudy (anonymous):

x=-8???

OpenStudy (austinl):

I believe so, I would have done it ratio style. -2×24=-48 -48÷6=-8 Therefore, x=-8

OpenStudy (anonymous):

ok i did it like the cross multiply. Thanks.